# On the loss of context-awareness in general instruction fine-tuning

### v2

@@ Abstract
Pre-trained Large Language Models (LLMs) require post-training methods such as supervised fine-tuning (SFT) on instruction-response pairs to enable instruction following. However, this process can potentially harm existing capabilities learned during pre-training. In this paper, we investigate the loss of context awareness after SFT, where context awareness is defined as the ability to extract and understand information from user-provided context and respond accordingly. We are the first to identify and show that the loss of context awareness, as reflected by the performance drop in the Needle-in-a-Haystack test, occurs in instruction fine-tuned LLMs when the chat template is applied to input prompts. We identify that the performance decline is partially caused by an attention bias toward different roles learned during conversational instruction fine-tuning. We validate our hypothesis by visualizing changes in attention allocation after the chat template is applied and manually steering the attention heads. Based on these observations, we propose a metric to select context-dependent examples from general instruction fine-tuning datasets. We then apply conditional instruction fine-tuning with a context-dependency indicator, enabling the model to learn context awareness from these selected examples. Empirical experiments on four context-dependent downstream tasks and three pre-trained LLMs of different sizes show that our method effectively mitigates the loss of context awareness without compromising general instruction-following capabilities. Given our findings, we strongly advocate for careful benchmarking of context awareness after instruction fine-tuning.
@@ 1 Introduction
Although some studies suggest that performance degradation can be mitigated or even eliminated through improved instruction fine-tuning methods (Longpre et al., 2023; Bai et al., 2022; Glaese et al., 2022), in this paper, we demonstrate that instruction fine-tuning specifically leads to the worsening of a model’s context awareness. Context awareness is defined as a model’s ability to accurately retrieve, process, and interpret specific information from user-provided context. Context awareness is highly relevant to the intrinsic hallucinations of LLMs (Huang et al., 2023) and crucial to the truthfulness of LLM-based chat models (Ouyang et al., 2022). It is also important for many real-world use cases, including retrieval augmented generalization (Khandelwal et al., 2020; Izacard et al., 2023; Xu et al., 2023b), in-context learning (Agarwal et al., 2024), and contextual question-answering (Rajpurkar et al., 2016; Choi et al., 2018; Dua et al., 2019).
@@
We demonstrate the loss of context awareness through evaluations of several popular, open-source LLMs using the Needle-in-a-Haystack (NIH) test (Kamradt, 2023), which requires a model to retrieve a given text “needle” from a long paragraph of irrelevant text. Performance on NIH tests deteriorates consistently after instruction fine-tuning across pretrained LLMs, regardless of their context window sizes, chat templates, architectures, or model sizes. Our analysis reveals that the decline in NIH performance is directly correlated with the application of chat templates during the instruction fine-tuning and inference phases. When these chat templates are not applied to instruction fine-tuned models, NIH performance not only recovers but in some instances surpasses that of their pretrained, non-fine-tuned counterparts.
@@
These observations suggest that the deterioration in NIH performance does not indicate a catastrophic loss of context-retrieval capabilities during instruction fine-tuning. Instead, the chat template appears to mask these underlying abilities by introducing systematic biases into the models’ behavior. Through analysis, we observe differences in attention allocation patterns in input tokens when comparing instruction fine-tuned models with and without applying chat templates. Specifically, we examine how the attention allocation shifts when tokens are marked as “user tokens” by the chat template. As illustrated in Figure 3, the application of chat templates leads to a redistribution of attention values: attention scores decrease for user input tokens while increasing for assistant tokens. We validate the relationship between attention score reallocation and context awareness through targeted intervention experiments. By manually increasing attention values assigned to user tokens, we partially restored the NIH performance. This confirms a direct connection between attention allocation and context processing capabilities.
@@
These findings motivate us to develop a fine-tuning strategy that mitigates attention bias acquired during instruction fine-tuning. Our approach stems from the intuition that the bias is learned from certain patterns in the training dataset. We develop a quantitative metric to assess the context-dependency of conversational instruction and response pairs based on attention allocation patterns. With this analysis, we discover that context-dependent examples are notably sparse in commonly used open-source instruction fine-tuning datasets. To address the limitation, we add an indicator token to the identified context-dependent instructions. After fine-tuning the model with this enhanced dataset, it learns to allocate increased attention to user tokens when the indicator is present.
@@
Our method is evaluated through experiments across three pretrained language models and four context-dependent tasks. Empirical results demonstrate that models fine-tuned using our method consistently achieve superior performance on context-dependent tasks compared to standard fine-tuning while maintaining similar performance on general tasks.
@@
  • We identify that the context awareness of LLMs deteriorates after supervised instruction fine-tuning with chat templates applied, as measured by the performance drop on the NIH task.
@@
  • We propose a quantitative metric to identify context-dependent instruction-response pairs from general instruction fine-tuning datasets.
  • We propose a fine-tuning strategy to insert a context-dependency indicator into identified instructions during instruction fine-tuning and mitigate the loss of context awareness by appending the indicator to instructions during inference.
@@ 2.1 Evaluating Context Awareness Through Needle-in-a-Haystack (NIH) Testing
In this section, we demonstrate the loss of context awareness after instruction fine-tuning with the needle-in-a-haystack (NIH) test.
@@
$\text{err}=\frac{1}{|K|}\sum_{w\in K}\mathbbm{1}(w\in output)$
where $K$ is the set of keywords in the targeted output and $output$ is the output of the LLM. We average the recall error across 400 NIH tests with different insertion locations and context lengths within the model’s context window. An example of the NIH prompt in our experiments is shown in Figure 1. When the chat template is applied to the prompt, the whole input prompt is partitioned into the user instruction input and model response, indicated by special role markers in the chat template (e.g., <|user|> and <|assistant|>). More details about the NIH tests can be found in Appendix A.2.
@@
We evaluate NIH on eight open-source language models from five model families with their corresponding chat templates. For each model, we compare the performance of the pretrained version (not instruction fine-tuned) and the official instruction-finetuned version released by the model provider. Here, we do not consider stronger closed-source models as their pretrained versions are unavailable. The context window lengths of these models range from $4K$ to $32K$.
@@
Figure 2: Average recall error (1 - recall) on NIH for different model series. We report the performance of official instruction-tuned models (both with and without chat templates) and their corresponding pretrained models from five model families, with sizes ranging from 7B to 27B. Note: Recall errors for some models are too small to be visible in the figure. Detailed numerical values can be found in Appendix B.1
The aforementioned phenomenon is consistent across models with different context window lengths, model families, chat templates, and model sizes.
@@ 2.2 Attention Allocation Bias Across Different Roles
Based on our observations, performance on NIH drops significantly when the chat template is applied. We hypothesize that the performance deterioration stems from the bias in instruction data and the bias is embedded in different roles marked by the chat template. When the model generates a response, it balances information from the input context and internal knowledge stored within its weights. It pays attention to user tokens to maintain consistency with the user-provided context while attending to previously generated assistant response tokens to maintain consistency with its output. If the model learns during fine-tuning that its internal knowledge is more important for generating high-quality responses, it may develop a bias that causes it to weigh user tokens less heavily. To verify our hypothesis, we visualize and compare the attention allocation between user and assistant tokens, both with and without chat templates.
@@ 2.3 Attention Steering to Compensate for Attention Bias
In the previous section, we observed a trend of decreased attention allocated to user tokens when chat templates were applied to instructional models, associated with a performance decline on the NIH task. A natural approach to mitigate this performance loss and further verify our hypothesis is to manually steer attention toward user tokens to compensate for the attention bias.
@@
where ${\mathbf{x}}$ and ${\mathbf{y}}$ are two tokens in the input sequence, $10<\alpha<$ 1, $U$ is the subset of all user tokens, and $Z$ is the normalization constant that renormalizes the altered attention scores across all tokens. $\text{Att}({\mathbf{x}},{\mathbf{y}})$ is the original attention weight from token ${\mathbf{x}}$ to token ${\mathbf{y}}$.
@@
Table 1: NIH and DROP performance with attention steering. DROP is a complicated reading comprehension task requiring discrete operations such as addition, sorting or counting on multiple pieces of information retrieved from the context. “Baseline” and “+ Attention Steering” are evaluated with chat templates. “w/o chat template” shows the NIH performance without the chat template for reference (same as the numbers reported in Figure 2).
Task Capabilities Model Name Baseline + Attention Steering w/o chat template
NIH sentence retrieval Llama-2-7B-Chat 0.9917 0.9935 0.9975
Llama-2-13B-Chat 0.9207 0.9375 0.9578
QuAC sentence retrieval Llama-2-7B-Chat 0.1330 0.1340 -
reading comprehension Llama-2-13B-Chat 0.1180 0.1230 -
DROP retrieval Llama-2-7B-Chat 0.4422 0.4416 -
math operation Llama-2-13B-Chat 0.4620 0.4584 -
We report the performance of attention steering in Table 1. Attention steering requires customized attention calculations for different heads and layers, which limits the use of several existing efficient attention implementations. Therefore, we are only able to apply attention steering to two Llama-2 models with 4,000-token context windows. In addition to the NIH test, we report the performance on two additional contextual QA tasks: QuAC (Choi et al., 2018) and DROP (Dua et al., 2019). Unlike NIH and QuAC, which retrieve exact sentences from the context, DROP requires the model not only to understand and retrieve relevant information from the context but also to apply discrete mathematical operations to the retrieved information. As shown in the table, attention steering can boost performance on simple context retrieval tasks such as NIH and QuAC. However, on DROP, which requires a more complex combination of different capabilities, performance with attention steering is negatively impacted. The deteriorating performance on DROP suggests that intervening in attention scores to emphasize user tokens, while improving context awareness, might impair other capabilities of the model.
In the next section, we introduce a finetuning strategy to mitigate the loss of context awareness with less side effects.
@@ 3 Instruction Fine-tuning with Context-dependency Indicators
In this section, we propose a fine-tuning strategy to mitigate the aforementioned loss of context awareness associated with chat roles. Our approach is based on the intuition that the model learns the bias from instruction-tuning datasets, which contain examples that rely less on user-provided context. Therefore, if we can identify “context-dependent” examples from these instruction fine-tuning datasets, we can apply several strategies — such as data augmentation, data selection, and conditional fine-tuning—to encourage the model to focus more on user-provided context.
@@ 3.1 Identifying context-dependent Instructions
To identify context-dependent instructions, we start by preparing a seed instruction fine-tuned model $M$, which can be the same or a weaker pretrained model fine-tuned with the original instruction fine-tuning dataset. We then define the context-dependency score for the $m^{\text{th}}$ turn response $Y_{m}$ given its instruction ${\mathbf{X}}_{m}$ and conversation history:
@@ 3.2 Instruction Fine-tuning with Context-dependency Indicators
A threshold $10<\beta<$ 1 can be selected after the context-dependency score is obtained for each instruction-response pair. A conversation turn $({\mathbf{X}}_{m},{\mathbf{Y}}_{m})$ with $s_{M}({\mathbf{Y}}_{m})>\beta$ is considered context-dependent. We append a special token [IND] to the user instruction ${\mathbf{X}}_{m}$ if it is context-dependent. In our implementation, the special token [IND] is added as an additional special token to the vocabulary to avoid conflicts with existing ones.
Our method is more general in terms of task and input format compared to related works, which synthesize context-dependent examples targeted toward specific tasks (e.g., contextual QA) (An et al., 2024). This makes our method more applicable to different types of user queries that require greater attention to user-provided context.
@@ 4.1.1 Models
We evaluate our method on three open-source pretrained large language models: TinyLlama-1.1B (Zhang et al., 2024a), Llama-2-7B (Touvron et al., 2023), and Llama-3-8B (Dubey et al., 2024). TinyLlama-1.1B is a 1.1B Llama model pretrained on 3 trillion tokens with a context window length of 2048. Llama-2-7B and Llama-3-8B have context windows of 4096 and 8192 tokens, respectively. Due to limited computational resources in academic labs, we are able to fine-tune only models with up to 8B parameters. We also truncate the training examples to 4096 tokens. We fine-tune Llama-2 and Llama-3 models using QLoRA (Dettmers et al., 2024) on transformer layers. Detailed hyperparameters can be found in Appendix A.1.1.

### v3

@@ Abstract
Pre-trained Large Language Models (LLMs) require post-training methods such as supervised fine-tuning (SFT) on instruction-response pairs to enable instruction following. However, this process can potentially harm existing capabilities learned during pre-training. In this paper, we investigate the loss of context awareness after SFT, where context awareness is defined as the ability to extract and understand information from user-provided context and respond accordingly. We identify and demonstrate that the loss of context awareness, particularly in open-source models, occurs in instruction fine-tuned LLMs when the chat template is applied to input prompts. We identify that the performance decline is associated with a bias toward different roles learned during conversational instruction fine-tuning. We demonstrate this correlation by visualizing changes in attention allocation after the chat template is applied and manually steering the attention heads. The bias can be learned from training examples that align with the model’s internal knowledge and rely less on the user-provided context to generate correct responses. Based on these observations, we propose a metric to identify context-dependent examples from general instruction fine-tuning datasets. We then apply conditional instruction fine-tuning with a context-dependency indicator, enabling the model to preserve context awareness after SFT. Empirical experiments on four context-dependent downstream tasks and three pre-trained LLMs of different sizes show that our method effectively mitigates the loss of context awareness without compromising general instruction-following capabilities.
@@ 1 Introduction
Although some studies suggest that performance degradation can be mitigated or even eliminated through improved instruction fine-tuning methods (Longpre et al., 2023; Bai et al., 2022; Glaese et al., 2022), in this paper, we demonstrate that instruction fine-tuning specifically leads to the worsening of a model’s context awareness in a series of open-source models. In this paper, context awareness is defined as a model’s ability to accurately retrieve, process, and interpret specific information from user-provided context. Context awareness is highly relevant to the intrinsic hallucinations of LLMs (Huang et al., 2023) and crucial to the truthfulness of LLM-based chat models (Ouyang et al., 2022). It is also important for many real-world use cases, including retrieval augmented generalization (Khandelwal et al., 2020; Izacard et al., 2023; Xu et al., 2023b), in-context learning (Agarwal et al., 2024), and contextual question-answering (Rajpurkar et al., 2016; Choi et al., 2018; Dua et al., 2019).
@@
We first demonstrate the loss of context awareness through evaluations of several popular, open-source LLMs using the Needle-in-a-Haystack (NIH) task. We show that while many pretrained models demonstrate near-perfect performance on NIH, their performance deteriorates consistently after SFT, regardless of context window sizes, chat templates, architectures, or model sizes. We show that this decline is correlated with the application of chat templates, which, however, are widely used and essential in building conversational LLM assistants. When these chat templates are removed from instruction fine-tuned models, NIH performance not only recovers but in some instances surpasses that of their pretrained, non-fine-tuned counterparts.
@@
These observations suggest that the deterioration in NIH performance does not indicate a catastrophic loss of context-retrieval capabilities during instruction fine-tuning. Instead, the chat template appears to mask these underlying abilities by introducing systematic biases into the models’ behavior. Through analysis, we observe differences in attention allocation patterns in input tokens when comparing instruction fine-tuned models with and without applying chat templates. Specifically, we examine how the attention allocation shifts when tokens are marked as “user tokens” by the chat template. As illustrated in Figure 3, the application of chat templates leads to a redistribution of attention values: attention scores decrease for user input tokens while increasing for assistant tokens. We further validate the relationship between attention score reallocation and context awareness through targeted intervention experiments. By manually increasing attention values assigned to user tokens, we partially restored the performance on simple context-relevant tasks.
@@
These findings motivate us to develop a fine-tuning strategy that mitigates attention bias acquired during instruction fine-tuning. Our approach stems from the intuition that the bias is learned from certain patterns in the training dataset, where for some examples the model does not need the context to generate correct answers. Therefore, we develop a quantitative metric to assess the context-dependency of conversational instruction and response pairs based on attention allocation patterns. We discover that context-dependent examples are notably sparse in commonly used open-source instruction fine-tuning datasets. To help the model distinguish examples with and without context-dependency during instruction finetuning, we add an indicator token to the identified context-dependent instructions. After fine-tuning the model with this enhanced dataset, it learns to allocate increased attention to user tokens when the indicator is present.
@@
We evaluate our method on three open-source pretrained language models and several context-dependent and general tasks. Empirical results demonstrate that models fine-tuned using our method consistently achieve superior performance on context-dependent tasks compared to standard fine-tuning while maintaining similar performance on general tasks.
@@
  • We identify that the context awareness of LLMs deteriorates after supervised instruction fine-tuning with chat templates applied, compared to pretrained models.
@@
  • We propose a quantitative metric to identify context-dependent instruction-response pairs from general instruction fine-tuning datasets. By inserting an indicator into identified instructions during SFT, we mitigate the loss of context awareness in instruct models when the indicator is added during inference, while preserving general performance.
@@ 2.1 Evaluating Context Awareness Through Needle-in-a-Haystack (NIH) Testing
Figure 2: Average recall error (1 - recall) on NIH for different model series. We report the performance of official instruction-tuned models (both with and without chat templates) and their corresponding pretrained models from five model families, with sizes ranging from 7B to 27B. Note: Errors on some models are too small to be visible in the figure. Detailed numerical values can be found in Appendix B.1
In this section, we demonstrate the loss of context awareness after instruction fine-tuning with the needle-in-a-haystack (NIH) test. We choose the NIH task because it relies less on instruction-following capabilities and provides a fairer comparison between pretrained models and instruct models in terms of context retrieval performance. We remove the newlines between context and needle in the original NIH test to increase difficulty and better discriminate among different models. An example of the NIH prompt used in our experiments is shown in Figure 1. We rerun the evaluation with different prompt templates for a more robust evaluation. More details can be found in Appendix A.2.
@@
$\text{recall}=\frac{1}{|K|}\sum_{w\in K}\mathbbm{1}(w\in output)$
$\text{err}=1-\text{recall}$
where $K$ is the set of keywords in the targeted output and $output$ is the output of the LLM. For all NIH evaluations, we calculate the recall on the first 100 generated tokens. We average the recall error across 400 NIH tests with different insertion locations and context lengths within the model’s context window. An example of the NIH prompt in our experiments is shown in Figure 1. When the chat template is applied to the prompt, the whole input prompt is partitioned into the user instruction input and model response, indicated by special role markers in the chat template (e.g., <|user|> and <|assistant|>). More details about the NIH tests can be found in Appendix A.2.
@@
We evaluate NIH on eight open-source language models from five model families. For each model, we compare the performance of the pretrained version (not instruction fine-tuned) and the official instruction-finetuned version released by the model provider. Here, we do not consider stronger closed-source models as their pretrained versions are unavailable. The context window lengths of these models range from $4K$ to $32K$.
@@
The aforementioned phenomenon is consistent across models with varying context window lengths, model families, chat templates, and small to medium model sizes. Although we are unable to conduct experiments on extremely large models, we believe that context awareness in medium-sized models remains important, considering their wide applications in cost-sensitive areas such as edge devices and small businesses.
@@ 2.2 Attention Allocation Bias Across Different Roles
Based on our observations, performance on NIH drops significantly when the chat template is applied. We hypothesize that the performance deterioration stems from the bias in instruction data and the bias is embedded in different roles marked by the chat template. When the model generates a response, it balances information from the input context and internal knowledge stored within its weights. It pays attention to user tokens to maintain consistency with the user-provided context while attending to previously generated assistant response tokens to maintain consistency with its output. If the model learns to assign lower importance to user-provided context and higher importance to its internal knowledge during SFT, it may develop a bias that causes it to weigh user tokens less. To provide evidence to our hypothesis, we visualize and compare the attention allocation between user and assistant tokens, both with and without chat templates.
@@ 2.3 Attention Steering to Compensate for Attention Bias
In the previous section, we observed a trend of decreased attention allocated to user tokens when chat templates were applied to instructional models, associated with a performance decline on the NIH task. A natural approach to further verify our hypothesis is to manually steer attention toward user tokens to compensate for the attention bias.
@@
where ${\mathbf{x}}$ and ${\mathbf{y}}$ are two tokens in the input sequence, $\alpha\in(0,1)$, $U$ is the subset of all user tokens, and $Z$ is the normalization constant that renormalizes the altered attention scores across all tokens. $\text{Att}({\mathbf{x}},{\mathbf{y}})$ is the original attention weight from token ${\mathbf{x}}$ to token ${\mathbf{y}}$.
@@
Table 1: NIH recall and QuAC/DROP containing score with attention steering. DROP is a complicated reading comprehension task requiring discrete operations such as addition, sorting or counting on multiple pieces of information retrieved from the context. “Baseline” and “+ Attention Steering” are evaluated with chat templates. “w/o chat template” shows the NIH performance without the chat template for reference (same as the numbers reported in Figure 2).
Task Capabilities Model Name Baseline + Attention Steering w/o chat template NIH sentence retrieval Llama-2-7B-Chat 0.9917 0.9935 0.9975 Llama-2-13B-Chat 0.9207 0.9375 0.9578 QuAC sentence retrieval Llama-2-7B-Chat 22.20 22.20 - reading comprehension Llama-2-13B-Chat 18.60 19.00 - DROP retrieval Llama-2-7B-Chat 44.22 44.16 - math operation Llama-2-13B-Chat 46.20 45.84 -
We report the performance of attention steering in Table 1. Attention steering requires customized attention calculations for different heads and layers, which limits the use of several existing efficient attention implementations. Therefore, we are only able to apply attention steering to two Llama-2 models with 4,000-token context windows. We report the recall on NIH task as well as the performances on two additional contextual QA tasks: QuAC (Choi et al., 2018) and DROP (Dua et al., 2019). Detailed descriptions and metrics of these tasks can be found in Section 4. Unlike NIH and QuAC, which retrieve exact sentences from the context, DROP requires the model not only to understand and retrieve relevant information from the context but also to apply discrete mathematical operations to the retrieved information. As shown in the table, attention steering can boost performance on simple context retrieval tasks such as NIH and QuAC. However, on DROP, which requires a more complex combination of different capabilities, performance with attention steering is negatively impacted. The deteriorating performance on DROP suggests that intervening in attention scores to emphasize user tokens, while improving context awareness, might impair other capabilities of the model.
Although we have shown the correlation between attention bias and loss of context awareness, we acknowledge that it may not fully explain the learned bias between user and assistant tokens. However, it can still serve as a proxy to identify how much the model relies on user-provided context to generate responses.
In the next section, we introduce a fine-tuning strategy to better mitigate the loss of context awareness with fewer side effects based on all of our aforementioned observations.
@@ 3 Instruction Fine-tuning with Context-dependency Indicators
Our method is based on the intuition that the model learns the bias from instruction-tuning datasets, which contain examples that rely less on user-provided context. Although one might anticipate that a model would learn to correctly decide whether to focus on context given the input prompt, the ”context-dependency” of a specific input prompt is not only prompt-dependent but also model-dependent. When the model has already learned the answer during pretraining, it does not need to rely on provided context to generate the correct answer. Therefore, if we can identify ”context-dependent” examples for each pretrained model and add an additional indicator to the prompts, we can mitigate the contextual attention bias learned during SFT, as the context-dependency factor has been included in the prompt.
@@ 3.1 Identifying context-dependent Instructions
We identify the context-dependent instructions by calculating the attention allocation on user tokens. We start by preparing a seed instruction fine-tuned model $M$, which can be the same or a weaker pretrained model fine-tuned with the original instruction fine-tuning dataset. We then define the context-dependency score for the $m^{\text{th}}$ turn response $Y_{m}$ given its instruction ${\mathbf{X}}_{m}$ and conversation history:
@@ 3.2 Instruction Fine-tuning with Context-dependency Indicators
A threshold $\beta\in(0,1)$ can be selected after the context-dependency score is obtained for each instruction-response pair. A conversation turn $({\mathbf{X}}_{m},{\mathbf{Y}}_{m})$ with $s_{M}({\mathbf{Y}}_{m})>\beta$ is considered context-dependent. We append a special token [IND] to the user instruction ${\mathbf{X}}_{m}$ if it is context-dependent. In our implementation, the special token [IND] is added as an additional special token to the vocabulary to avoid conflicts with existing ones.
@@ 4.1.1 Models
We evaluate our method on three open-source pretrained large language models: TinyLlama-1.1B (Zhang et al., 2024a), Llama-2-7B (Touvron et al., 2023), and Llama-3-8B (Dubey et al., 2024). TinyLlama-1.1B is a 1.1B Llama model pretrained on 3 trillion tokens with a context window length of 2048. Llama-2-7B and Llama-3-8B have context windows of 4096 and 8192 tokens, respectively. Due to limited computational resources in academic labs, we are able to fine-tune only models with up to 8B parameters. We also truncate the training examples to 4096 tokens. Detailed hyperparameters can be found in Appendix A.1.1.